# Trajectory World Models for Heterogeneous Environments

### v1

@@ Abstract
Heterogeneity in sensors and actuators across environments poses a significant challenge to building large-scale pre-trained world models on top of this low-dimensional sensor information. In this work, we explore pre-training world models for heterogeneous environments by addressing key transfer barriers in both data diversity and model flexibility. We introduce UniTraj, a unified dataset comprising over one million trajectories from 80 environments, designed to scale data while preserving critical diversity. Additionally, we propose TrajWorld, a novel architecture capable of flexibly handling varying sensor and actuator information and capturing environment dynamics in-context. Pre-training TrajWorld on UniTraj demonstrates significant improvements in transition prediction and achieves a new state-of-the-art for off-policy evaluation. To the best of our knowledge, this work, for the first time, demonstrates the transfer benefits of world models across heterogeneous and complex control environments.
@@ 1 Introduction
In recent years, machine learning has been revolutionized by foundation models pre-trained on large-scale, diverse data (Achiam et al., 2023; Oquab et al., 2024; Kirillov et al., 2023). General world models have also been realized through pre-training, enabled by the homogeneity present within massive and diverse datasets of specific modalities, such as text (Wang et al., 2024b; Gu et al., 2024), images (Zhou et al., 2024), and videos (Seo et al., 2022; Wu et al., 2024a, b; Agarwal et al., 2025). However, a unique challenge of world models from Internet AI is commonly overlooked or circumvented: the heterogeneity inherent in sensor and actuator information, which means proprioceptive data, such as joint positions and velocities, as well as optional target positions, vary significantly across environments. Failing to properly address this heterogeneity can result in no transfer, or even negative transfer.
@@
Figure 2: Illustration of pre-training a world model from heterogeneous environments, with each environment labeled by its state and action dimensions. A Trajectory World Model, designed for flexibility in handling divergent state and action definitions, demonstrates effective positive transfer across distinct, heterogeneous and complex control environments.
@@
By pre-training our flexible TrajWorld architecture on the diverse and massive UniTraj dataset, we demonstrate, for the first time, the transfer benefits of world models across heterogeneous and complex control environments. Fine-tuning TrajWorld on 15 datasets from three previously unseen environments (Fu et al., 2020) significantly reduces transition prediction errors for both in-distribution and out-of-distribution actions (as shown in Figure 1). This improved predictive accuracy also translates to our state-of-the-art performance on off-policy evaluation (OPE) tasks (Fu et al., 2021), enabling the offline evaluation and selection of a set of complex policies for best performance.
@@
  • For the first time, our experiments demonstrate positive world model transfer across diverse and complex environments, achieving significant improvements in both transition prediction and policy evaluation.
@@ 4.3 Towards a General Trajectory World Model
5 Experiments
In this section, we test the following hypotheses:
  • Large-scale trajectory pre-training can generalize effectively and even enable zero-shot generalization, contrary to the common belief. (Section 5.1)
  • TrajWorld outperforms alternative architectures for transition prediction when transferring dynamics knowledge to new environments. (Section 5.2)
@@ 5 Experiments
  • TrajWorld leverages the general dynamics knowledge acquired from pre-training to improve performance in downstream tasks. (Section 5.3)
@@ 5.1 Zero-shot Generalization
We further find that TrajWorld, when trained on the large-scale UniTraj dataset, is also capable of zero-shot generalizing to unseen environments, Cart-2-Pole and Cart-3-Pole from DMC (Figure 4(b) and 10). Specifically, TrajWorld successfully infers the influence of the action value (pushing force) on the state dimension (cart position) and accurately predicts the outcomes for different action sequences performed subsequently.
@@ 5.2 Transition Prediction
We compare our approach against two baselines: an ensemble of MLPs (Chua et al., 2018), widely adopted for dynamics modeling, and TDM (Schubert et al., 2023), which is similar to our model but uses one-dimensional attention. Each baseline is evaluated both for training from scratch and fine-tuning pre-trained ones on the same UniTraj dataset as TrajWorld. To enable pre-training, we pad the state and action vectors with zeros to match the same dimensionality for MLP. Additionally, we include our model trained from scratch for comparison.
@@ Results.
Figure 1 presents the aggregated mean absolute error of 75 train-test dataset pairs for various models. TrajWorld outperforms all baselines, highlighting the effectiveness of its pre-training strategy and architecture design. Notably, MLP Ensemble with pre-training performs worse than its non-pre-trained counterpart, emphasizing the importance of careful model design for world modeling across heterogeneous environments. While TDM also benefits significantly from pre-training, it still lags behind TrajWorld. This is likely because TDM naively treats everything as a 1D sequence, neglecting the unique problem structures. In contrast, TrajWorld explicitly models variate relationships and temporal transitions, leveraging different facets of dynamics knowledge from the pre-training. Moreover, TDM predicts variants sequentially, which may accumulate errors and lead to less accurate results.
In Figure 5, we further show detailed prediction error results for TrajWorld compared to its non-pre-trained counterparts. In 12 out of 15 training datasets, fine-tuned TrajWorld achieves a lower average prediction error across 5 test datasets, further validating the effectiveness of pre-training. Moreover, the transfer benefits are evident in both in-distribution and out-of-distribution scenarios.
@@ 5.3 Off-Policy Evaluation
Off-policy evaluation (OPE) estimates the value of a target policy using an offline transition dataset collected by a separate behavior policy. It is commonly used to select the most performant policy from a set of candidates when online evaluation is too costly to be practical. This task provides an ideal evaluation scenario for world models, as value estimation can be acquired by rolling out the target policy within the learned world model.
@@ 5.4 Analysis
Effects of pre-training dataset.
To further investigate which aspects contribute to the pre-training benefits, we train TrajWorld on a modified UniTraj dataset with the Modular RL and TD-MPC2 data removed. Despite this modification, TrajWorld still demonstrates significant pre-training advantages (Appendix C.6). This highlights that the benefits of pre-training are derived from the diversity of the entire UniTraj dataset, rather than relying exclusively on data from domains that are relatively similar to target environments.
@@ 6 Related Work
Zero-padding to fit a maximum length (Yu et al., 2020a; Hansen et al., 2024; Schmied et al., 2024; Seo et al., 2022) or using separate neural network heads (Wang et al., 2024a; D’Eramo et al., 2020) hinders knowledge transfer between heterogeneous environments with mismatched or differently sized state and action spaces. Previous work has resorted to flexible architectures like graph neural networks (Huang et al., 2020; Kurin et al., 2021) and Transformers (Gupta et al., 2022; Hong et al., 2021) for policy learning. Our method leverages a similar architecture for world modeling (Janner et al., 2021; Zhang et al., 2021), but with a novel two-dimensional attention design to enhance cross-environment transfer.
@@ B.4.3 Metrics
B.5 Computational Cost
@@ C.1 Detailed Prediction Error for Baselines
We report the prediction error for MLP Ensemble and TDM in Figure 8 and 9, respectively.
@@ C.2 Quantitative Results for Off-Policy Evaluation
C.4 Additional Zero-shot Cross-Environment Transfer
@@
C.5 Additional Variate Attention Visualization
We present the variate attention maps of our TrajWorld model across all six layers, comparing a fine-tuned model and a model trained from scratch, in Figures 11 and 12.
@@ C.5 Additional Variate Attention Visualization
C.6 Ablation Study on Pre-training Dataset

### v2

@@ Abstract
Heterogeneity in sensors and actuators across environments poses a significant challenge to building large-scale pre-trained world models on top of this low-dimensional sensor information. In this work, we explore pre-training world models for heterogeneous environments by addressing key transfer barriers in both data diversity and model flexibility. We introduce UniTraj, a unified dataset comprising over one million trajectories from 80 environments, designed to scale data while preserving critical diversity. Additionally, we propose TrajWorld, a novel architecture capable of flexibly handling varying sensor and actuator information and capturing environment dynamics in-context. Pre-training TrajWorld on UniTraj yields substantial gains in transition prediction, achieves a new state-of-the-art for off-policy evaluation, and also delivers superior online performance of model predictive control. To the best of our knowledge, this work, for the first time, demonstrates the transfer benefits of world models across heterogeneous and complex control environments. Code and data are available at https://github.com/thuml/TrajWorld.
@@ 1 Introduction
In recent years, machine learning has been revolutionized by foundation models pre-trained on large-scale, diverse data (Achiam et al., 2023; Oquab et al., 2024; Kirillov et al., 2023). General world models have also been realized through pre-training, enabled by the homogeneity present within massive and diverse datasets of specific modalities, such as text (Wang et al., 2024b; Gu et al., 2024; Chae et al., 2025; Wu et al., 2025), images (Zhou et al., 2024), and videos (Seo et al., 2022; Wu et al., 2024a; Bruce et al., 2024; Wu et al., 2024b; Agarwal et al., 2025). However, a unique challenge of world models from Internet AI is commonly overlooked or circumvented: the heterogeneity inherent in sensor and actuator information, which means proprioceptive data, such as joint positions and velocities, as well as optional target positions, vary significantly across environments. Failing to properly address this heterogeneity can result in no transfer or even negative transfer.
@@
Figure 2: Illustration of pre-training a world model from heterogeneous environments, with each environment labeled by its state and action dimensions. A Trajectory World Model, designed for flexibility in handling divergent state and action definitions, demonstrates effective positive transfer across distinct, heterogeneous, and complex control environments.
@@
By pre-training our flexible TrajWorld architecture on the diverse and massive UniTraj dataset, we demonstrate, for the first time, the transfer benefits of world models across heterogeneous and complex control environments. Fine-tuning TrajWorld on 15 datasets from three previously unseen environments (Fu et al., 2020) significantly reduces transition prediction errors for both in-distribution and out-of-distribution actions (as shown in Figure 1). This improved predictive accuracy also translates to our state-of-the-art performance on off-policy evaluation (OPE) tasks (Fu et al., 2021), enabling the offline evaluation and selection of a set of complex policies for best performance. Furthermore, it also manifests in superior online performance with model predictive control (MPC).
@@
  • For the first time, our experiments demonstrate positive world model transfer across diverse and complex environments, resulting in simultaneous and significant improvements in transition prediction, off-policy policy, and model predictive control.
@@ 4.3 Towards a General Trajectory World Model
5 Experiments
In this section, we test the following hypotheses:
  • Large-scale trajectory pre-training can generalize effectively and even enable zero-shot generalization, contrary to the common belief (Section 5.1).
  • TrajWorld outperforms alternative architectures for transition prediction when transferring dynamics knowledge to new environments (Section 5.2).
  • TrajWorld leverages the general dynamics knowledge acquired from pre-training to improve performance in downstream tasks (Section 5.3).
@@ 5.1 Zero-shot Generalization
We further find that TrajWorld, when trained on the large-scale UniTraj dataset, is also capable of zero-shot generalizing to unseen environments, Cart-2-Pole and Cart-3-Pole from DMC (Figure 4(b) and 13). Specifically, TrajWorld successfully infers the influence of the action value (pushing force) on the state dimension (cart position) and accurately predicts the outcomes for different action sequences performed subsequently.
@@ 5.2 Transition Prediction
We compare our approach against two baselines: an ensemble of MLPs (Chua et al., 2018), widely adopted for dynamics modeling, and TDM (Schubert et al., 2023), which is similar to our model but flattens inputs and uses one-dimensional attention. Each baseline is evaluated both for training from scratch and fine-tuning pre-trained ones on the same UniTraj dataset as TrajWorld. To enable pre-training, we pad the state and action vectors with zeros to match the same dimensionality for MLP. Additionally, we compare with our model trained from scratch.
@@ Results.
Figure 1 presents the aggregated mean absolute error of 75 train-test dataset pairs for various models. TrajWorld outperforms all baselines, highlighting the effectiveness of its pre-training strategy and architecture design. Notably, MLP Ensemble with pre-training performs worse than its non-pre-trained counterpart, emphasizing the importance of careful model design for world modeling across heterogeneous environments. While TDM also benefits significantly from pre-training, it still lags behind TrajWorld. This is likely because TDM naively treats everything as a 1D sequence, neglecting the unique problem structures. In contrast, TrajWorld explicitly models variate relationships and temporal transitions, leveraging different facets of dynamics knowledge from the pre-training. Moreover, TDM predicts variants sequentially, which may accumulate errors and lead to less accurate results, whereas TrajWorld predicts all variables jointly, mitigating compounding errors.
In Figure 5, we further show detailed prediction error results for TrajWorld compared to its non-pre-trained counterparts. In 12 out of 15 training datasets, fine-tuned TrajWorld achieves a lower average prediction error across 5 test datasets, further validating the effectiveness of pre-training. Moreover, the transfer benefits are evident in both in-distribution and out-of-distribution scenarios, indicating that the model generalizes well even when trained and tested on transitions collected from different policies.
@@ 5.3 Off-Policy Evaluation
Off-policy evaluation (OPE) estimates the value of a target policy using an offline transition dataset collected by a separate behavior policy. It is commonly used to select the most performant policy from a set of candidates when online evaluation is too costly to be practical. This task provides an ideal evaluation scenario for world models, as value estimation can be acquired by rolling out the target policy within the learned world model. This is particularly advantageous for evaluating long-horizon predictions, where direct environment interaction is infeasible and model accuracy over extended timeframes is critical.
@@ Results.
Figure 7: Model predictive control (MPC) results with proposal policies across three environments, averaged over three random seeds.
5.4 Model Predictive Control
Model predictive control (MPC) selects actions by optimizing predicted future rewards over a finite horizon using a learned world model, making it well-suited for evaluating world model performance in online control settings.
Setup.
We evaluate MPC performance in a practical scenario where world models trained on medium-replay datasets are used to enhance medium-level proposal policies through model predictive control. Specifically, we utilize three medium-replay datasets from D4RL and medium-level policies from DOPE. Additionally, we experiment with MPC using a random shooting planner. Implementation details are provided in Appendix B.5.
Baselines.
As in the previous section, we compare our method against the MLP Ensemble and TDM baselines, evaluting both both from-scratch and fine-tuned variants.
Results.
Figure 7 presents results for MPC with proposal policies. Overall, MPC using TrajWorld yields the highest-performing agents, outperforming both baseline models and its from-scratch counterpart. We find that MPC leads to significant gains in the Hopper and Walker2D environments, but has limited effects in HalfCheetah, likely due to its inherent stability and lower risk of failure. In contrast, Hopper and Walker2D are fragile, and our world models help prevent unsafe actions, leading to better planning. Notably, the TDM model exhibits negative transfer in the MPC with proposal policies setting, despite showing positive transfer in transition prediction and off-policy evaluation.
In the random shooting setting, the planner’s limited ability to sample high-quality actions hinders its effective utilization of model predictions, leading to consistently poor performance across all world models. Nevertheless, TrajWorld demonstrates comparatively better results under these limitations. See Appendix C.4 for detailed results.
@@ 5.5 Analysis
(a) Transition prediction error.
(b) Model predictive control.
Figure 9: Effects of pre-training scale and diversity. (a) Aggregated transition prediction error. (b) Model predictive control performance on Walker2D and Hopper. All results are obtained from models fine-tuned from pre-trained TrajWorld on different subsets of UniTraj.
@@
Effects of pre-training scale and diversity
We assess the impact of dataset scale and diversity by pre-training three TrajWorld variants on subsets of UniTraj: a 1/10 sample, a 1/100 sample, and the JAT subset (purely expert trajectories from five environments), followed by fine-tuning on downstream environments. For transition prediction, we adopt a challenging setup where models are fine-tuned on expert data per environment but evaluated across all data levels. Model-predictive control (MPC) results are also reported. As shown in Figure 9, all subset-pretrained models outperform training from scratch, but fall short of the model trained on the full UniTraj dataset. This reveals a desirable scaling trend: larger and more diverse pre-training data consistently lead to better generalization. These findings highlight the value of heterogeneous pre-training on large-scale datasets.
@@ 6 Related Work
Zero-padding to fit a maximum length (Yu et al., 2020a; Hansen et al., 2024; Schmied et al., 2024; Seo et al., 2022) or using separate neural network heads (Wang et al., 2024a; D’Eramo et al., 2020) hinders knowledge transfer between heterogeneous environments with mismatched or differently sized state and action spaces. Previous work has resorted to flexible architectures like graph neural networks (Huang et al., 2020; Kurin et al., 2021) and Transformers (Gupta et al., 2022; Hong et al., 2021) for policy learning. Our method leverages a similar architecture for world modeling (Janner et al., 2021; Zhang et al., 2021), but introduces the two-dimensional attention design for the first time in this context. More importantly, no for computational efficiency, as in prior work of other fields (Ho et al., 2019; Arnab et al., 2021; Nayakanti et al., 2023), we show its benefits for cross-environment transfer.
@@ 7 Conclusion
Acknowledgements
This work was supported by the National Natural Science Foundation of China (U2342217 and 62021002), the BNRist Innovation Fund (BNR2024RC01010), and the National Engineering Research Center for Big Data Software.
@@ A.2 List of Environments
A.3 Sampling Weights
We manually weighted different subsets, trying to balance size and diversity. The sample weights are shown in Table 3.
Subsets ExoRL RLU JAT DB-1 TD-MPC2 Modular RL
(Unnormalized) sampling weight 75 5 90 1 90 30
Table 3: Sampling weights of subsets for pre-training with UniTraj dataset.
@@ B.4.3 Metrics
B.5 Model Predictive Control
We evaluate model predictive control (MPC) performance under two planning settings: policy proposal and random shooting.
Policy proposal setting.
Action candidates are generated by perturbing the output of a learned action policy with Gaussian noise. Specifically, we first query the policy to obtain a mean action sequence and then add zero-mean Gaussian noise to each action in the sequence. This results in a set of diverse trajectories centered around the policy’s behavior.
Random shooting setting.
Candidate action sequences are sampled directly from a Gaussian distribution without guidance from a learned policy. Each trajectory is independently sampled by drawing actions from a zero-mean Gaussian distribution with a fixed standard deviation.
Hyperparameters.
For both settings, we use a sample size of 128 candidate trajectories per MPC rollout, across all environments. The best-performing action sequence is selected based on predicted cumulative reward computed using the world model.
The planning horizon is set based on the characteristics of each environment. Specifically, we use a horizon of 25 steps for both HalfCheetah and Walker2D, while a longer horizon of 50 steps is adopted for Hopper. This extended horizon for Hopper helps mitigate short-sighted planning behavior, which is particularly detrimental in this more fragile environment.
To ensure optimal performance across different world models, the standard deviation of the Gaussian noise used for action sampling is tuned individually for each environment. The noise level is set to 0.05 for Hopper, 0.2 for Walker2D, and 0.025 for HalfCheetah. These values were empirically selected to balance exploration and stability during trajectory sampling.
These settings are used consistently in all experiments involving MPC in this work. The same configurations are applied for evaluating all world models, ensuring fair comparison.
B.6 Computational Cost
@@ C.1 Detailed Prediction Error for Baselines
We report the prediction error for MLP Ensemble and TDM in Figure 10 and 11, respectively.
@@ C.2 Quantitative Results for Off-Policy Evaluation
C.4 Additional Model Predictive Control Results
Env. MLP (w/o PT) MLP (w/ PT) TDM (w/o PT) TDM (w/ PT) TW (w/o PT) TW (w/ PT) Proposal
Hopper 948$\pm$61 1091$\pm$125 1287$\pm$26 1117$\pm$145 1090$\pm$225 1401$\pm$236 1078$\pm$143
Walker 3353$\pm$83 3465$\pm$20 3056$\pm$236 2619$\pm$36 2422$\pm$455 3427$\pm$370 3049$\pm$104
HalfCheetah 5645$\pm$10 5692$\pm$19 5611$\pm$85 5647$\pm$25 5858$\pm$17 5809$\pm$15 5697$\pm$30
Table 8: Quantitative results of all model-based methods (TW=TrajWorld) for MPC with action proposal, averaged over 3 seeds.
Quantitative results with proposal policies.
We report quantitative results on MPC with action proposal in Table 8.
MPC with random shooting planner.
Figure 12 presents MPC results using a random shooting planner with models trained on different datasets.
Computational efficiecny.
TrajWorld predicts all variates jointly, unlike TDM which processes them sequentially. This leads to a major speedup: MPC for 1000 environment steps in HalfCheetah takes 40 minutes with TDM, but only 3 minutes with TrajWorld.
Figure 12: Model predictive control (MPC) results using a random shooting planner, averaged across three random seeds. The proposal policy line indicates the performance of a random action-sampling strategy.
C.5 Additional Zero-shot Cross-Environment Transfer
@@
Comparison with baselines.
We also provide zero-shot prediction from other baselines in Figure 14. As shown, in an unseen environment, both TDM and MLP baselines fail to generalize, producing incorrect predictions and failing to capture the underlying state-action relationship at all. Specifically, TDM fails to predict how push forces from two opposite directions lead to different x positions. On the other hand, MLP fails to produce any reasonable results with extreme error accumulation.
Figure 14: Zero-shot predictions from different pre-trained models on two Cart-2-Pole trajectories that share the same 10 context steps but diverge thereafter due to different future actions.
Cart-3-pole environment.
@@
C.6 Additional Variate Attention Visualization
We present the variate attention maps of our TrajWorld model across all six layers, comparing a fine-tuned model and a model trained from scratch, in Figures 15 and 16.
@@ C.6 Additional Variate Attention Visualization
C.7 Additional Ablation Study on Pre-training Dataset
@@ Appendix D Extended Discussion
Discussion with Schubert et al. (2023).
We demonstrate positive transfer to complex downstream environments such as Walker2D, not only for offline transition prediction and policy evaluation, but also for online MPC, which Schubert et al. (2023) did not. Our work differentiate from theirs in: (1) Setting: Instead of fine-tuning with $10^{4}$ episodes for MPC with random shooting, we more practically fine-tune with $10^{2}$ episodes for MPC with proposal policies; (2) Data diversity: Our UniTraj dataset emphasizes distribution diversity, rather than using pure expert trajectories; (3) Architecture: TrajWorld incorporates inductive biases tailored to the 2D structure of trajectory data for enhanced transferability. Notably, TDM exhibits negative transfer in our practical MPC setting. We believe our work complements and extends Schubert et al. (2023), offering new insights to the community.